OpenStudy (anonymous):

If h(x)=integral (x^2 to 2) sqr root (1+t^4)dt

zepdrix (zepdrix):

\[\Large\rm h(x)=\int\limits_{x^2}^2 \sqrt{1+t^4}~dt\]Do we need to find h'(x)?

OpenStudy (anonymous):

my guess, which certainly could be wrong, is that you are asked for the derivative of this thing like i said, i could be wrong

OpenStudy (anonymous):

@zepdrix guessed it too

zepdrix (zepdrix):

heh :p

OpenStudy (anonymous):

Sorry, I forgot to ask the question. I am looking for h'(x).

OpenStudy (anonymous):

\[\Large h(x)=\int\limits_{x^2}^2 \sqrt{1+t^4}~dt\] \[=\Large h(x)=-\int\limits_{2}^{x^2}\sqrt{1+t^4}~dt\] since the derivative of the integral is the integrand, replace \(t\) by \(x^2\) and then via the chain rule, multiply the result by \(2x\)